OpenStudy (babyslapmafro):

How do you express the following decimals as fractions? .9999... 5.373737... thank you

OpenStudy (tkhunny):

The usual trick is a simple multiplication. I'll do the first one. You do the second. N = 0.99999.... Then 10N = 9.9999999.... 10N - N = 9.99999... - 0.99999... 9N = 9 N = 1 <== Weren't expecting that, were you? :-) Clearly, there is some interesting conversation to be had around the multiplication and addition of infinite decimals, but that is the idea. M = 5.3737.... Go!

OpenStudy (babyslapmafro):

ok thanks

OpenStudy (anonymous):

\[\huge\frac{ 532 }{ 99 }\]

OpenStudy (anonymous):

second one!!

OpenStudy (babyslapmafro):

Do I follow the same steps as the first one? I can't get a decent fraction for N

OpenStudy (babyslapmafro):

N = 5.373737.... Then 10N = 53.73737.... 10N - N = 48.36363... 9N = 48.36363

OpenStudy (anonymous):

there is a shortcut for this question!!

OpenStudy (tkhunny):

ahahah! Define "decent". M = 5.373737... What I didn't say, is the unmotivated multiplication needs to be enough for the entire repetitive part of the decimals. When I had 0.99999, I used 10, because only a single digit repeats. You'll need 100! 100M = 537.373737....

OpenStudy (anonymous):

u dont need to do all these steps!!

OpenStudy (babyslapmafro):

oh i see

OpenStudy (tkhunny):

It is important to note that unique answers don't care how you find them. There may be shortcuts and alternate methodologies. You need to figure out what makes the most sense to you and helps you along your way.

OpenStudy (babyslapmafro):

do i also do 100M-M or is it 100M-10M

OpenStudy (tkhunny):

There was no 10M. M = 5.373737... 100M = 537.373737... 100M - M = ??

OpenStudy (babyslapmafro):

ok thank i got 532/9 now

OpenStudy (tkhunny):

I hope not. 100M - M = 99M

OpenStudy (babyslapmafro):

ya thats what i meant to type

OpenStudy (tkhunny):

You should see quickly that the number of repeating digits is always the number of 9s you get in the denominator.

OpenStudy (babyslapmafro):

529/99

OpenStudy (babyslapmafro):

AHHH 532/99

OpenStudy (tkhunny):

Care to take a few more tries? :-)

OpenStudy (babyslapmafro):

sure